Hi-Tek Manufacturing, Inc., makes two types of industrial component parts—the B300 and the T500. An absorption costing incomeRequired: 1. Compute the product margins for the B300 and T500 under the companys traditional costing system. 2. Compute theCompute the product margins ToP the B300 and 1500 Under the companys tradltlonal costlng system. (Round your intermediate caCompute the product margins for B300 and T500 under the activity-based costing system. (Negative product margins should be inPrepare a quantitative comparison of the traditional and activity-based cost assignments. (Round your intermediate calculatioB300 T500 Total % of % of Amount Amount Amount Traditional Cost System Total cost assigned to products $ $ Total cost $ 0 Tot
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Hi-Tek
Answer 1
Calculation of manufacturing overhead per unit under traditional costing B300 T500 Total
Direct Labor cost (D)      120,100.00      42,900.00       163,000.00
Manufacturing overhead (A)       500,700.00
Manufacturing overhead per direct labor cost (C=A/D)                   3.07
Allocated Manufacturing overhead (E=C*D)      368,920.67 131,779.33       500,700.00
B300 T500 Total
Selling Price (B)                20.00             40.00
No. of units sold (F)        60,200.00      12,800.00         73,000.00
Sales (G=B*F) 1,204,000.00 512,000.00    1,716,000.00
Direct Materials      400,900.00 162,700.00       563,600.00
Direct Labor      120,100.00      42,900.00       163,000.00
Manufacturing overhead      368,920.67 131,779.33       500,700.00
Gross Margin      314,079.33 174,620.67       488,700.00
Advertising Expenses        54,000.00 101,000.00       155,000.00
Net Margin (H)      260,079.33     73,620.67       333,700.00
Product Margin (I=H/F)                  4.32                5.75
Answer 2 J K L=J/K M N=M*L O P=L*O Q=M+O R=N+P
Calculation of product margin under ABC costing: B300 T500 Total
Activities Overhead cost Total Activity Activity Rate Activity Amount Activity Amount Activity Amount
Machining (Machine Hours)      209,610.00 153,000.00                   1.37 90,100.00       123,437.00 62,900.00        86,173.00 153,000.00        209,610.00
Setups (Setups Hours)      130,290.00           303.00               430.00          73.00         31,390.00        230.00        98,900.00          303.00        130,290.00
Product Sustaining (number of products)      100,200.00                2.00         50,100.00             1.00         50,100.00             1.00        50,100.00               2.00        100,200.00
Total Overhead      440,100.00       204,927.00      235,173.00        440,100.00
Direct Materials       400,900.00      162,700.00        563,600.00
Direct Labor       120,100.00        42,900.00        163,000.00
Advertising Expenses         54,000.00      101,000.00        155,000.00
Total Cost       779,927.00      541,773.00     1,321,700.00
Sales    1,204,000.00      512,000.00     1,716,000.00
Margin (S)       424,073.00       (29,773.00)        394,300.00
Product Margin (T=S/F)                   7.04                (2.33)
Answer 3 B300 T500 Total
Traditional costing Amount % Amount % Amount %
Direct Materials      400,900.00 45.05%       162,700.00 48.22%       563,600.00 45.92%
Direct Labor      120,100.00 13.50%         42,900.00 12.72%       163,000.00 13.28%
Manufacturing overhead      368,920.67 41.46%       131,779.33 39.06%       500,700.00 40.80%
Total cost assigned to products      889,920.67       337,379.33 1,227,300.00
Advertising Expenses       155,000.00
Total cost 1,382,300.00
Answer 3 B300 T500 Total
ABC costing Amount % Amount % Amount %
Direct Costs
Direct Materials      400,900.00 51.40%       162,700.00 30.03%       563,600.00 42.64%
Direct Labor      120,100.00 15.40%         42,900.00 7.92%       163,000.00 12.33%
Advertising Expenses        54,000.00 6.92%       101,000.00 18.64%       155,000.00 11.73%
Indirect Costs
Machining (Machine Hours)      123,437.00 15.83%         86,173.00 15.91%       209,610.00 15.86%
Setups (Setups Hours)        31,390.00 4.02%         98,900.00 18.25%       130,290.00 9.86%
Product Sustaining (number of products)        50,100.00 6.42%         50,100.00 9.25%       100,200.00 7.58%
Total cost assigned to products      779,927.00       541,773.00 1,321,700.00
Costs not assigned to products
Organizational sustaining costs         60,600.00
Total cost 1,382,300.00